OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 7, -11, and 2 + 8i

hero (hero):

Hint: Multiply f(x) = (x - 7)(x + 11)(x - (2 + 8i))

OpenStudy (anonymous):

Um I am a bit unsure, I mulitpled the first part and got (x^2+4x-77) but how do I multiply w (x - (2 + 8i)?

hero (hero):

(x - (2 + 8i)) = (x - 2 - 8i)

OpenStudy (anonymous):

(154+616 i)-(85+32 i) x-79 x^2+x^3 ?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

but these are my choices: f(x) = x4 - 9x3 - 56x2 + 290x - 5236 f(x) = x4 - 9x3 + 56x2 - 290x + 5236 f(x) = x4 - 145x2 + 580x - 5236 f(x) = x4 - 25x2 + 580x - 5236

hero (hero):

Try this: f(x) = (x - 7)(x + 11)(x - (2 + 8i))(x + (2 + 8i))

OpenStudy (anonymous):

yeah because on wolfram, what you are giving me has the same zeros as my problem...so idk..

hero (hero):

Well, that's the clue then

hero (hero):

Use wolfram alpha to figure out which one of your answer choices also has the same zeroes.

hero (hero):

Wolframalpha calculated the zeroes of the polynomial we came up with: http://www.wolframalpha.com/input/?i=solve+%28x+-+7%29%28x+%2B+11%29%28x+-+%282+%2B+8i%29%29%28x+%2B+%282+%2B+8i%29%29+%3D+0

hero (hero):

If you use wolframalpha in the same manner for the other possible answer choices, you should be able to find the polynomial that has the same zeroes. Once you find it, that will be the correct choice.

OpenStudy (anonymous):

ended up f(x) = x4 - 25x2 + 580x - 5236 thanks for your help :) finished my work lol

hero (hero):

I just realized that I wrote the original one down wrong. It should have been: f(x) = (x - 7)(x + 11)(x - 2 + 8i)(x - 2 - 8i) = 0

hero (hero):

^If you multiply that out, you get x^4 - 25x^2 + 580x - 5236

OpenStudy (anonymous):

oh lol. well you were better than me anyway :p